# Driving as a Diagnostic Tool: Scenario-based Cognitive Assessment in Older Drivers From Driving Video

### v1

@@ Abstract
We introduce scenario-based cognitive status identification in older drivers from Naturalistic driving videos and large vision models. In recent times, cognitive decline, including Alzheimer’s disease (AD) and mild cognitive impairment (MCI), is often underdiagnosed due to the time-consuming and costly nature of current diagnostic methods. By analyzing real-world driving behavior captured through in-vehicle systems, this research aims to extract “digital fingerprints” that correlate with functional decline and clinical features of MCI and AD. Moreover, modern large vision models can draw meaningful insights from everyday driving patterns of older patients to early detect cognitive decline. We propose a framework that uses large vision models and naturalistic driving videos to analyze driver behavior, classify cognitive status and predict disease progression. We leverage the strong relationship between real-world driving behavior as an observation of the current cognitive status of the drivers where the vehicle can be utilized as a “diagnostic tool”. Our method identifies early warning signs of functional impairment, contributing to proactive intervention strategies. This work enhances early detection and supports the development of scalable, non-invasive monitoring systems to mitigate the growing societal and economic burden of cognitive decline in the aging population.
@@
Naturalistic driving, Large Vision model, Cognitive decline, driving scenario
@@
[label1]organization=Department of Electrical & Computer Engineering,addressline=Iowa State University, city=Ames, state=IA, country=USA
@@ 1 Introduction
Cognitive decline in the aging population is a critical health concern that manifests across a spectrum of stages. The pre-dementia stage, commonly referred to as mild cognitive impairment (MCI), is characterized by measurable cognitive impairment with minimal or no functional changes in daily living. Individuals at this stage, often in their 70s, may exhibit early signs of cognitive decline without severe disruptions to independence. As the condition progresses to dementia, particularly Alzheimer’s disease (AD), more severe cognitive and functional impairments emerge, ultimately diminishing the ability to live independently. While MCI is not exclusive to AD, it often serves as an early indicator of the disease and highlights the need for early detection and intervention.
@@
This study focuses on individuals exhibiting MCI due to AD, as determined by clinical characteristics. The timely identification of such individuals provides an opportunity to slow disease progression, improve quality of life, and reduce the societal burden associated with advanced stages of dementia. Traditional methods for diagnosing cognitive decline are often time-consuming, costly, and reliant on clinical visits or complex neuroimaging techniques, which may not be practical for widespread use [1, 2]. To address this, our research explores an innovative approach using real-world driving behavior as a passive diagnostic tool.
Prior research on cognitive decline has explored a variety of methods to detect early changes in individuals at risk of AD. Neuroimaging techniques, such as magnetic resonance imaging (MRI) [3] and positron emission tomography (PET), have been widely used to identify structural and functional brain changes associated with MCI and AD. These methods are often paired with biomarkers, such as amyloid-beta and tau proteins, to improve diagnostic accuracy. However, the high cost and limited accessibility of these techniques hinder their widespread application.
Wearable sensor technologies [4, 5] have also gained traction in recent years. Studies have demonstrated the utility of devices that monitor gait patterns, heart rate variability, and sleep quality to assess functional and cognitive impairments. Machine learning models [6, 7] applied to sensor data have shown promise in predicting the progression of cognitive decline. However, these approaches often require participant compliance and are limited to specific contexts, such as controlled environments or clinical trials.
Traditional methods like road tests and driving simulators are commonly used to assess fitness-to-drive [8] but are limited to controlled conditions and face challenges related to objectivity, availability, and cost-effectiveness. To address these limitations, driving research has shifted towards naturalistic assessments using Sensor devices [9] installed in participants’ personal vehicles, enabling the collection of real-world driving behavior data.
Previous studies [10, 11] have demonstrated that everyday driving behavior, such as reduced driving distances, fewer unique destinations, and smaller driving spaces, is associated with symptomatic Alzheimer’s disease (AD). However, limited research has investigated the impact of preclinical AD on driving behavior [12], highlighting a critical gap in understanding early-stage cognitive decline.
In contrast, our approach capitalizes on passive monitoring of real-world driving behavior, which offers a non-invasive, cost-effective, and scalable solution. By combining driving video data with advanced computer vision techniques, we aim to bridge the gap between traditional clinical diagnostics and real-world applications, providing an innovative pathway for early detection of cognitive decline.
@@ 2.1 RWRAD Dataset
This dataset was collected in collaboration with the University of Nebraska Medical Center and encompasses comprehensive neuropsychological assessments, clinical tests, and lab results to investigate cognitive impairment and its progression.
Table 1: Study participant characteristics
@@
Weekly Driving Range: 5 - 300 miles (8 - 483 km) 85.9 miles (138.2 km) 75.2 miles (121.0 km)
@@
Participants were classified using established criteria for mild cognitive impairment (MCI), which identifies individuals with cognitive changes that do not yet produce major functional impairments. Table 1 shows the detailed demographic and cognitive layout of the participants. Key classification methods include:
  • 1.
    Jak criteria: [13] Two tests in one domain scoring more than 1 standard deviation (SD) below the mean, offering higher longitudinal accuracy but detecting fewer MCI cases.
  • 2.
    Peterson criteria: [14] One test in one domain scoring more than 1.5 SD below the mean, detecting more MCI cases but with reduced longitudinal accuracy.
MCI classifications primarily identify amnestic impairments, often associated with short-term and episodic memory, though impairments in other domains, such as language and visuospatial skills (atypical presentations), are increasingly recognized. As cognitive impairment progresses, it often becomes a multi-domain process, ultimately impacting functional abilities and transitioning to the dementia stage of clinical Alzheimer’s disease (AD). The study design for collecting in-lab data focused on assessing functional abilities and activities through driving behavior in various scenarios. Key questions included identifying participants’ characteristics and health status, determining cognitive impairments and their probable link to AD, evaluating co-morbid sources of cognitive decline, and assessing overall health, cognitive function, mobility, vision, and signs of dependence, depression or quality-of-life decline.
To address challenges associated with compliance and self-reporting accuracy in cognitively impaired participants, a study partner was recruited for each participant. Data collection included contributions from both participants and study partners, with comparative analyses used to ensure data reliability and robustness. This longitudinal dataset provides a valuable resource for studying the progression of cognitive decline and functional impairment.
@@
Clinical evaluation of MCI due to AD is primarily a rule-out process aimed at determining whether the primary impairment is due to another type of dementia. However, the majority of people with AD ($\sim$60%) also have other types of dementia [15]. While this study does not exclude individuals with co-morbid non-AD dementias, clinical assessments are designed to capture these data. Participants recruited into this study are included because they exhibit primary AD. It is important to note that AD presents a complex and heterogeneous landscape, making its study both challenging and critical.
In addition to clinical diagnosis and functional assessments, composite cognitive scores can offer a quantitative summary of neuropsychological performance across multiple domains. A composite cognition score (COGSTAT) integrates standardized measures from several well-established neuropsychological tests. This approach captures impairments in memory, language, visuospatial skills, and executive function—core domains often affected in AD and MCI. The formula for COGSTAT is as follows:
$\text{COGSTAT}=350$ $+\frac{C-38.7}{11.1}+\frac{L-26.2}{3.5}+\frac{R-10.08}{3.2}$ (1)
$+\frac{B-4.4}{2.4}+\frac{R_{cs}-31}{3.7}+\frac{R_{rs}-15.7}{5.7}+\frac{T-46.1}{32.6}$
Where:
  • 1.
    $C$ = Controlled Oral Word Association test (continuous value)
  • 2.
    $L$ = Judgment of Line Orientation test
  • 3.
    $R$ = Auditory Verbal Learning test: 30-minute recall score
  • 4.
    $B$ = Number of errors in Benton Visual Retention test
  • 5.
    $R_{cs}$ = Rey-O Complex Figure test: copy score
  • 6.
    $R_{rs}$ = Rey-O Complex Figure test: 30-minute recall score
  • 7.
    $T$ = Trail Making test Part B: time in seconds
While such neuropsychological testing provides a comprehensive assessment of cognitive abilities, it often requires in-person clinical visits, participant compliance, and substantial time and resources — all of which may be challenging for individuals in the early stages of AD. In this context, analyzing naturalistic driving video offers a promising, non-invasive alternative to capture subtle behavioral cues of cognitive decline in real-world settings. Studies show that behavioral patterns in everyday driving can be an alternative and effective approach to detect early-stage AD, complementing traditional cognitive testing and potentially enabling earlier intervention.
2.2 Video Data Pre-pocessing
Figure 1: Omaha most transited route database GIS map (Omaha, NE)
Figure 2: Breakdown of missing and redundant video in RWRAD dataset
@@ 2.2 Video Data Pre-pocessing
Figure 3: Cognitive score (a) COGSTAT, (b) MoCA distribution across subjects ($N=69$). We combined the AD and MCI subjects into a unified class“AD-aging” and performed two-class classification with the “Normal-aging”.
@@
Figure 4: Preparation of video samples for specific scenarios. The top row represents the rollout of all driving videos and drives across subjects. These video files from different subject blocks are filtered based on scenarios and used as input to the Large Vision model.
@@
Figure 5: Overview of the video-to-embedding generation. Given, input video segments $\mathcal{V}=\{V_{k}^{(D_{i},Sc_{j})}\}\mid i=1,\dots,d;\;j=1,\dots,k;\;k=1,\dots,n\}$ from different scenarios $Sc_{j}\in\mathcal{S}$, a vision model $F_{\theta}$ maps each $v_{i}$ to an embedding $Z_{i}=F_{\theta}(V_{i})\in\mathbb{R}^{m}$. The embedding space is analyzed to assess inter-group separation using a distance metric. Scenario $Sc_{1}$ exhibits greater class-wise separation than $Sc_{2}$, i.e., $\text{Dist}_{Sc_{1}}>\text{Dist}_{Sc_{2}}$, indicating higher discriminative potential for identifying cognitive states. Selected embeddings are passed to a classifier $H(Z_{i})\rightarrow\hat{Y}_{i}$ for downstream cognitive classification.
Data was collected during a 3-month baseline - including driving video, actigraphy, sleep log, driving behavior and in-lab data by the University of Nebraska Medical Center in Omaha. Ninety legally licensed and active, aged between 65 - 90 years old participants were recruited with a study partner near Omaha, NE. They were recruited through community outreach, including flyers, local news, and community engagement [16]. Over a 2-year period, driving data were collected in 3-month intervals using custom-installed “Black Boxes” in each participant’s personal vehicle. These systems passively recorded driving behavior throughout each trip, capturing synchronized forward roadway and cabin video alongside sensor data from GPS, speed, accelerometer, gyroscope, and throttle. All sensor data were processed at a 1Hz sampling rate. All participants provided informed consent in accordance with institutional protocols.
@@
While analyzing the collected driving video data, we encountered some missing and corrupted video clips that required filtering before analysis. Fig. 2 presents the breakdown of the processed driving data for our analysis. While processing the in-vehicle Blackbox video data, we identified a total of 520,662.7 minutes of driving footage, segmented into pure drive duration (72.6%), blackframe drive duration (4.1%), and missing drive duration (23.3%). For downstream analysis, only valid segments—comprising both pure and black-frame video—were retained. Segments with missing or corrupted data were excluded to ensure input quality and consistency for driver behavior modeling and cognitive health assessment tasks.
@@
Subjects included in this study were required to pass at least a predefined fraction of the Omaha most transited route road segments during the three-month baseline period. This criterion was applied to ensure adequate video data coverage for each participant and to make sure we have diverse driving scenario exposure in the dataset. By focusing on these criteria, our framework extracts meaningful insights into the cognitive status of aging drivers, correlating observed behaviors with functional decline. For our analysis, a total of $N=69$ subjects were utilized for video-based analysis using Year-1 data, after excluding the subjects without class labels and drivers outside the Omaha region. Figure 3 shows the distribution of COGSTAT scores across three aging groups— AD, MCI, and Normal aging. Among 69 subjects, 35 drivers (26 with MCI and 9 with Alzheimer’s Disease) were categorized as “AD-aging”, while 34 were classified as “Normal-aging”. This balanced distribution enabled a robust comparison between normal and AD aging groups in terms of cognitive test scores and driving behavior. Subjects with Alzheimer’s consistently show the lowest scores, while normal-aging individuals cluster at the higher end. MCI subjects span the intermediate range, overlapping with both groups. It also highlights the transitional nature of cognitive decline and the inherent classification challenge.
@@
To ensure consistency and relevance in our analysis, we focused on short segments. We selected roadway segments where a drive can be $\leq 0.31$ mile or $\leq 500$ meters and defined it as a “route segment” (discussed in 3.1). These segments provide concentrated snapshots of driver behavior under controlled road conditions. This approach minimizes variability due to changes in driving context or prolonged periods of observation, which may dilute critical behavioral signals. Additionally, the time of day and traffic conditions (peak vs. off-peak hours) were considered to account for environmental stressors that could uniquely challenge drivers with Alzheimer’s clinical syndrome (ACS).
2.2.1 Scenario-specific Video Sample

### v2

@@ Abstract
We introduce scenario-based cognitive status identification in older drivers from naturalistic driving videos, leveraging large vision models. In recent times, cognitive decline including Dementia and Mild Cognitive Impairment (MCI), is often underdiagnosed due to the time-consuming and costly nature of current diagnostic methods. By analyzing real-world driving behavior captured through in-vehicle sensors, this study aims to extract “digital fingerprints” that correlate with functional decline and clinical features of dementia. Moreover, modern large vision models can draw meaningful insights from everyday driving patterns across different roadway scenarios to early detect cognitive decline. We propose a framework that uses large vision models and naturalistic driving videos to analyze driver behavior, identify cognitive status and predict disease progression. We leverage the strong relationship between real-world driving behavior as an observation of the current cognitive status of the drivers where the vehicle can be utilized as a “diagnostic tool”. Our method identifies early warning signs of functional impairment, contributing to proactive intervention strategies. This work enhances early detection and supports the development of scalable, non-invasive monitoring systems to mitigate the growing societal and economic burden of cognitive decline in the aging population.
@@
Naturalistic driving, Vision model, Cognitive decline, Driving scenario
journal: anonymous
@@
[label1]organization=Department of Electrical & Computer Engineering,addressline=Iowa State University, city=Ames, state=IA, postcode=50011, country=USA
@@ 1 Introduction
Cognitive decline in the aging population is a critical health concern that manifests across a spectrum of stages. The pre-dementia stage, commonly referred to as mild cognitive impairment (MCI), is characterized by measurable cognitive impairment with minimal or no functional changes in daily living [1, 2]. Individuals at this stage (often in their 70s) may exhibit early signs of cognitive decline without severe disruptions to independence. As the condition progresses to dementia, more severe cognitive and functional impairments emerge, ultimately diminishing the ability to live independently [3]. While MCI is not exclusive to dementia, it often serves as an early indicator of the disease and highlights the need for early detection and intervention [4].
@@
The timely identification of MCI provides an opportunity to slow disease progression, improve quality of life, and reduce the societal burden associated with advanced stages of dementia. Traditional methods for diagnosing cognitive decline are often time-consuming, costly, and reliant on clinical visits or complex neuroimaging techniques, which may not be practical for widespread use [2, 5, 6]. To address this, our research explores an innovative approach using real-world driving behavior as a passive diagnostic tool [7, 8]. This study focuses on individuals exhibiting MCI due to mild dementia, as determined by clinical characteristics [9].
Prior research on cognitive decline has explored a variety of methods to detect early changes in individuals at risk of dementia. Neuroimaging techniques, such as magnetic resonance imaging (MRI) [10, 11] and positron emission tomography (PET) [12, 13] have been widely used to identify structural and functional brain changes associated with dementia. These methods are often paired with biomarkers, such as amyloid-beta and tau proteins [14], blood-based biomarkers [15] to improve diagnostic accuracy. However, the high cost and limited accessibility of these techniques hinder their widespread application [16]. Wearable sensor technologies [17, 18] have also gained traction in recent years. Studies have demonstrated the utility of devices that monitor gait patterns, heart rate variability, and sleep quality to assess functional and cognitive impairments. Machine learning models [19, 20, 21] applied to sensor data have shown promise in predicting the progression of cognitive decline. However, these approaches often require participant compliance and are limited to specific contexts, such as controlled environments or clinical trials [22]. Some studies [23, 24, 25] demonstrated that everyday driving behavior, such as reduced driving distances, fewer unique destinations, and smaller driving spaces, is associated with symptomatic MCI. However, limited research has investigated the impact of preclinical MCI on driving behavior [26, 27], highlighting a critical gap in understanding early-stage cognitive decline. Recently, driving research has shifted towards naturalistic assessments using sensor devices installed in participants’ personal vehicles, enabling the collection of real-world driving behavior data [7, 26].
While neuropsychological testing provides a comprehensive assessment of cognitive abilities, it often requires in-person clinical visits, participant compliance, and substantial time and resources — all of which may be challenging for individuals in the early stages of MCI [28]. In this context, our approach capitalizes on passive monitoring of real-world driving behavior, which offers a non-invasive, cost-effective, and scalable solution. Analyzing naturalistic driving data offers a promising alternative to capture subtle behavioral cues of cognitive decline in real-world settings [29, 30]. Studies show that behavioral patterns in everyday driving can be an alternative and effective approach to detect early-stage MCI, complementing traditional cognitive testing and potentially enabling earlier intervention [29, 31, 26]. By combining driving video data with advanced computer vision techniques, we aim to discover the relationship between driving behavior and cognitive assessment, providing a digital “diagnostic tool” for early detection of cognitive decline. We propose a scenario-based framework that identifies discriminative behavioral cues and shows how scenario-specific analysis can reveal early cognitive decline more effectively than aggregate driving metrics.
@@ 2.1 RWRAD Dataset
The dataset was collected in collaboration with the University of Nebraska Medical Center (UNMC) and includes comprehensive neuropsychological assessments, clinical tests and naturalistic driving data aimed at investigating cognitive impairment and its progression among older adults.
2.1.1 Participants
These data were derived from a cohort of community-dwelling older adults (aged 65 - 90 years) enrolled in longitudinal studies on dementia and driving behavior, conducted at the UNMC Mind and Brain Health Labs between March 2021 and January 2024. Recruitment procedures followed prior work [32, 29], involving community outreach efforts (e.g., flyers, social media, local news, discussions at senior organizations) and pre-screening to ensure eligibility based on defined inclusion criteria (see A). All participants lived independently in Nebraska or neighboring states (e.g., Iowa), held a valid driver’s license, were proficient in English and provided informed consent. Exclusion criteria were applied to minimize potential confounding effects from underlying health conditions and are detailed in A. The study protocol was approved by the UNMC Institutional Review Board (IRB 0522-20-FB) and all participants provided written informed consent prior to participation. A summary overview of the participants is shown in Table 1.
2.1.2 Data collection
Data collection includes recording driving video, sensor signals, actigraphy, sleep log, driving behavior and in-lab cognitive assessment conducted at UNMC. Over a 2-year period, driving data were collected in 3-month intervals using custom-installed “Black Box” systems in each participant’s personal vehicle. The system continuously recorded driving behavior throughout each trip over a 12-week period. It captured synchronized video (front roadway and cabin), vehicle sensor data (GPS, accelerometer, gyroscope), and on-board diagnostics data (speed, rpm, engine throttle). All data were recorded at a 1Hz sampling rate from engine start to stop (ignition cycle). Each ignition cycle indicates a distinct “drive” for the corresponding driver, excluding trips with zero mileage or speed<5<5mph (e.g., parking lot or driveway maneuvers) or if participants reported periods of non-driving (e.g., vacations, illness).
Table 1: Summary characteristics of the participants in RWRAD dataset
@@ 2.1.2 Data collection
Weekly Driving (miles) Range: 5 - 300 85.9 75.2
COGSTAT Score Range: 300 - 800 643.22 106.52
@@
2.1.3 Cognitive Status
Cognitive status was determined using the 2018 NIA-AA research criteria [9]. To assess cognitive status, participants completed a series of tests across five domains – memory, language, visuospatial abilities, executive function and attention (see B). Specialist clinicians in dementia and Alzheimer’s reviewed the neuropsychological evaluations for cognition and daily functioning from the National Alzheimer’s Coordinating Center (NACC) Unified Data Set (UDS) version 3 [33]. They reached a consensus on cognitive staging based on participants performance across the five cognitive domains. Raw test scores were converted into standardized $Z$-scores, adjusted for age, sex, and education. These $Z$-scores were then used to define cognitive impairment. As cognitive impairment progresses, it often becomes a multi-domain process, ultimately impacting functional abilities and transitioning to the dementia stage. The in-lab data collection focused on assessing functional abilities and activities through driving behavior in various roadway scenarios. Cognitive performance was evaluated using two composite cognitive scores. First, the COGSTAT score was utilized as a measure of cognitive function [34]. It was derived from 14 neuropsychological tests evenly distributed across those five cognitive domains. For each test, an age-adjusted $Z$-score was calculated and transformed into a $T$-score (mean = 50, SD = 10) using the formula $T=10(Z\times 10+50)$. The composite scores required complete data without missing tests to ensure validity. Second, the Montreal Cognitive Assessment (MoCA) score was used as a clinical screening tool designed to detect clinically significant cognitive change and age-related cognitive decline [35]. Both the MoCA and the full neuropsychological battery were administered to all participants. To address challenges associated with compliance and self-reporting accuracy in cognitively impaired participants, a study partner was recruited for each participant. After two weeks, participants returned to the study site with their partner for evaluations of cognitive status and daily functioning.
@@ 2.1.3 Cognitive Status
Participants were classified based on the presence of both cognitive and functional decline. Those without any impairment in cognition or daily functioning were labeled as cognitively unimpaired (CU) or “Normal-aging”. Individuals who demonstrated cognitive decline but maintained functional independence were classified as having mild cognitive impairment (MCI). When cognitive decline was accompanied by functional impairment, participants were considered to have dementia. Given the small number of dementia cases ($N=9$), participants with MCI and dementia were combined into a single category termed Mild Dementia (MD) or “MD-aging”, similar to stage 4 in [9]. Clinical evaluation of MCI is primarily a rule-out process aimed at determining whether the primary impairment is due to another type of dementia. However, the majority of people with Alzheimer’s ($\sim$60%) also have other types of dementia [36]. While this study does not exclude individuals with co-morbid dementias, clinical assessments were designed to capture these data.
2.2 Video Data Processing
Figure 1: Breakdown of the driving data in the RWRAD dataset
@@ 2.2 Video Data Processing
Figure 2: Cognitive score (a) COGSTAT, (b) MoCA distribution across subjects ($N=69$). We combined the MD and MCI subjects into a unified class“MD-aging” and performed two-class classification with the “Normal-aging”.
Figure 3: GIS map of the Omaha most transited route database (Omaha, NE)
@@
Figure 4: Scenario-specific video sampling. The top row represents the rollout of all valid drives and driving videos across subjects in the Omaha most transited route. From each subject $D_{i}$, valid drives $dr_{1},dr_{2},\dots,dr_{n}$} and their associated video samples $V_{1},V_{2},\dots,V_{m}$} were mapped into scenario-specific subsets. These video samples were aggregated and used as input to the large vision model for driving pattern analysis.
@@
Figure 5: Overview of the large vision model-based framework. Given, the input video segments $\mathcal{V}=\{V_{k}^{(D_{i},Sc_{j})}\}$ where $i=1,\dots,d;\;j=1,\dots,k;\;k=1,\dots,n$ from different scenarios $Sc_{j}\in\mathcal{S}$, a vision model $F_{\theta}$ mapped each $V_{i}$ to a vector space $Z_{i}=F_{\theta}(V_{i})\in\mathbb{R}^{m}$. The vector space was analyzed to assess inter-group separation using a distance metric. Scenario $Sc_{1}$ showed significant class-wise separation than $Sc_{2}$, i.e., $\text{Dist}_{Sc_{1}}>\text{Dist}_{Sc_{2}}$, indicating higher discriminative potential for distinguishing the cognitive groups.
@@
Quality driving video is essential for reliable video-based modeling since missing or corrupted data can hurt the validity of findings. Ensuring rigorous processing of the collected video clips was therefore a crucial step before conducting our analysis. While processing the collected driving video data, we encountered some missing and corrupted video clips that required filtering and formatting. Figure 1 presents the breakdown of the processed driving data for the RWRAD dataset. The in-vehicle Blackbox system recorded a total of 520,662.7 minutes of driving data, segmented into valid drive duration (73.7%), Blackframe drive duration (3.9%) and Missing drive duration (22.4%). For the video-based analysis, only valid driving data was retained. Drives with missing or corrupted data and Black-frame data were excluded to ensure input data quality and consistency for driver behavior modeling.
@@
Subjects with their corresponding driving videos were filtered out for only those drives that occurred on the Omaha most transited route, a predefined set of road segments derived from the RWRAD dataset (see Figure 3). This criterion was applied to ensure adequate video data coverage for each participant and to make sure we have diverse driving scenario exposure in the video samples. In addition, the common driving route ensures that individual drivers commute on a shared set of road segments, reducing variability due to route differences. This allows the model to focus on relevant driving patterns rather than route features. For our analysis, a total of $N=69$ subjects were filtered out, after excluding the subjects without cognitive status labels and drivers outside the Omaha region. Figure 2 shows the distribution of the cognitive scores across three aging groups— MD, MCI, and Normal. Among 69 subjects, 35 drivers (26 MCI and 9 MD combined) were categorized as “MD-aging”, while 34 were classified as “Normal-aging”. This balanced distribution enabled a robust comparison between normal and MD-aging groups in terms of cognitive test scores and driving behavior. The MD-aging subjects consistently show the lowest cognitive scores, while normal-aging individuals cluster at the higher end. MCI subjects span the intermediate range, overlapping with both groups. It also highlights the transitional nature of cognitive decline and the inherent classification challenge.
@@
To ensure consistency and relevance in our analysis, we focused on short driving segments. We selected roadway segments where the drive distance can be $\leq 0.31$ mile or $\leq 500$ meters and defined it as a “route segment” (discussed in 3.1). These segments provide concentrated snapshots of driver behavior under controlled road conditions. This approach minimizes variability due to changes in driving context or prolonged periods of noisy observation. Additionally, the time of day and traffic conditions (peak vs. off-peak hours) were considered to account for environmental stressors that could uniquely challenge drivers with mild dementia. Next, the scenario-specific video sampling method is described in section 2.3.
2.3 Scenario-specific Video Sampling